Evaluate the collision frequency of H2 molecules in a gas at 1.00atm and 25 degrees celsius
Answer #1

According to the kinetic Molecular Theory,

=

Where, = Collision frequency

= Root-mean-square velocity of the molecules

= Mean free path

Root-mean-square velocity :

The formula relating the rms velocity to the temperature and molar mass is:

vrms =

Where, R = The universal Gas constant

T = The temperature

M =Tthe molar mass

For H2 at 25 °C ,

T = ( 25 + 273.15 ) K = 298.15 K

M = 2.016 g.mol-1 = 2.016 10-3 kg.mol-1

vrms = =

vrms = m.s-1 = 1920 m.s-1

The mean free path :

For a hydrogen molecule, Kinetic diameter σ = 289 pm.

The formula for the mean free path is

λ =

Where = Mean free path

R = 8.314 10-5 bar.m3.K-1 .mol-1

T = 298.15 K

NA = Avogadro number = 6.022 1023 mol-1

P = Pressure = 1.00 atm = 1.01325 bar

λ =

=

= m

= 1.094 10-7 m = 10.94 nm

Now Collision frequency,

=

=

= 1.755 10-10 s-1

There are 1.755 10-10 collisions per second.